an aeroplane travelling at a speed of 500 km h titles at an angle of 30° as it makes a turn . what is the radius of curve

Grade:12th pass

1 Answers

Saurav
5 years ago

Are you asking a ncert textbook question .
Ok,
 velocity of the plane (v)=500kmph=139 mps.
Angle at which it made the turn(x) = 30°
Let' s take acceleration due to gravity(g) to be 9.8mps^-2.
Then ,

Using the formula,
v^2/rg = tanx.
where r= radius of the loop.
Putting the values we get
19321/9.8r=0.57

r= 19321/9.8×0.57
r=3458.8 metres.

Or ,the radius of the loop executed by the plane is 3.45 km.